What weight of NaCl in 800g acetone is needed to bring boiling point to 60°C? (Boiling

point of pure acetone= 55.95°C ; Kb of acetone = 1.71 °C/m)


1
Expert's answer
2022-03-10T03:20:02-0500

∆t = i×Kb×m

∆t = (60-55.95) = 4.05

Kb = 1.71°C/m

m = 800g = 0.8kg

i = 1

4.05 = (1)(1.71)(x/0.8)

4.05 = 2.1375x

x = 1.891 mol

Molar mass of NaCl= 58.44g/mol

(58.44×1.891) = 110.45

= 110.45g